### v1

# Topological complexity of Milnor fibration

Abstract.
In this paper we discover a connection between the Milnor fibration theory and current research trends in topological robotics. The configuration space and workspace are often described as subspaces of some Euclidean spaces. The work map is a continuous map which assigns to each state of the configuration space the position of the end-effector at that state. The tasking planning problem consists in study the complexity and design of algorithms controlling the task performed by the robot manipulator. We study the tasking planning problem for the Milnor fibration of analytic function-germs in the real and complex cases. The tasking algorithms strongly depends of the Milnor fibration theorem.
Key words and phrases:
Milnor fibration, cross-section, sectional number, LS category, robotics, algorithms.
2010 Mathematics Subject Classification:
Primary 32S55, 68T40; Secondary 52C35,70Q05
The second author would like to thank grant #2022/03270-8 and grant#2022/16695-7, São Paulo Research Foundation (FAPESP) for financial support.
1. Introduction
In this paper “space” means a topological space, and fibrations are taken in the Hurewicz sense.
From robotics (see [3]), the configuration space $C$ of a mechanical system $\mathcal{S}$ is defined as the space of all possible states of $\mathcal{S}$. The robot workspace $W$ consists of all points that can be reached by the robot end-effector. The configuration space and workspace are often described as a subspace of some Euclidean space $\mathbb{R}^{n}$ and $\mathbb{R}^{p}$, respectively. The work map or kinematic map is a continuous map
$f:C\to W$
which assigns to each state of the configuration space the position of the end-effector at that state. For instance, we consider the robot arm as shown in Figure 1. In this case the configuration space is the torus $S^{1}\times S^{1}$ and the workspace is the $2$-sphere $S^{2}$, and the work map is given by $f(\alpha,\beta)=\left(\cos{\alpha}\cos{\beta},\cos{\alpha}\sin{\beta},\sin{\alpha}\right)$, see [12].
Figure 1. Robot arm (RR).
The planning problem consists in implementing algorithms (we will say tasking planning algorithms) controlling the task performed by the robot manipulator. The input of such an algorithm, see Pavešić [12], are pairs $(a,A)\in C\times W$ of points of the configuration space and workspace, respectively. The output for such a pair ought to be a continuous path $\alpha$ in the configuration space $C$ such that $\alpha(0)=a$ and $f(\alpha(1))=A$, see Figure 2.
Figure 2. The planning problem. The input of an algorithm are pairs $(a,A)\in C\times W$ and the output for such a pair ought to be a path in the configuration space $\alpha\in PC$ such that $\alpha(0)=a$ and $f(\alpha(1))=A$.
For practical purposes, a tasking planning algorithm should depend continuously on the pair of points $(a,A)$. Unfortunately, if a (global) continuous tasking planning algorithm of $f:C\to W$ exists implies that $f$ admits a global continuous cross-section and $W$ is contractible (see Lemma 3.8). Yet, if $f$ does not admit a global continuous cross-section or $W$ is not contractible, we could care about finding tasking planning algorithms $s$ defined only on a certain open set of $C\times W$, to which we refer as the domain of definition of $s$. In these terms, a tasking planner of $f$ is a set of local continuous tasking planning algorithms whose domains of definition cover $C\times W$.
The topological complexity of $f$, TC$(f)$, is then the minimal cardinality among tasking planners of $f$ (see Definition 3.9), while a tasking planner of $f$ is said to be optimal if its cardinality is TC$(f)$. In the case that $f$ does not admit a tasking planner with a finite number of domains of definition, we write $\text{TC}(f)=\infty$. The design of explicit motion planners that are reasonably close to optimal is one of the challenges of modern robotics (see, for example Latombe [8] and LaValle [9]).
Remark 1.1.
It is known that singularities of the kinematic map implies a reduction in freedom of movement in the working space that arises in certain joint configurations (see [5], [12]) and thus the complexity of motion planning can be higher. Thus, it is reasonable to ask whether it is possible to lower the topological complexity of work maps by restricting them in a neighbourhood of singularities.
In the book [11] Singular Points of Complex Hypersurfaces, John W. Milnor studied singularities of complex hypersurfaces and showed how to attach a locally trivial smooth fiber bundle to each singular point in order to extract topological data from it. Milnor proved the existence of such fibration structures associated to holomorphic functions and also to real analytic maps, but in the latter case for isolated critical points. Both fibrations are currently known as the Milnor fibrations. The real Milnor fibrations have been extended in several directions for isolated and not isolated singularities. Some interesting results toward that can be found, for example, in [1], [2], [10] and [14].
We consider $n>p\geq 2$ and $f:(\mathbb{R}^{n},0)\longrightarrow(\mathbb{R}^{p},0)$ be an analytic map germ which satisfies the Milnor’s conditions $(a)$ and $(b)$ (see Definition 2.3). Instead to study the topological complexity of $f$ (which can be higher, see Remark 1.1) we can consider the Milnor fibration as work map:
$f_{\mid}:f^{-1}(S^{p-1}_{\eta})\cap B^{n}_{\epsilon}\to S^{p-1}_{\eta},$
where $0<\eta\ll\epsilon\leq\epsilon_{0},$ and $\epsilon_{0}$ is a Milnor’s radius for $f$ at origin (see Corollary 2.7). Recall that, $E=E(\eta,\epsilon)=f^{-1}(S^{p-1}_{\eta})\cap B^{n}_{\epsilon}$ denotes the Milnor tube. We will show that it is possible to lower the topological complexity of work maps by restricting them in a neighbourhood of singularities (see Item (1) from Remark 4.1), so there is a greater need for work to understand it. The purpose of this work is to study the topological complexity of the Milnor fibration.
The paper is organized as follows: We start with a brief review about the existence of Milnor fibration and its topology in the complex and real cases (Section 2). In addition, we present Remark 2.8. Proposition 2.12 presents a characterisation of the existence of cross-sections for Milnor fibrations in terms of homotopy groups of the fiber. In Section 3 we review the notion of topological complexity. We present and prove Lemma 3.2, its proof is very technical and will be use in the last section. Proposition 3.5 presents a characterisation of the connectivity of the Milnor fiber in terms of the sectional number. Proposition 3.7 presents a necessary condition for the connectivity of the Milnor fiber in terms of the singular cohomology of the Milnor tube. Lemma 3.8 presents a motivation for the definition of topological complexity. The algorithms presented in Example 3.11 are keys to our purpose. Another fundamental result is Proposition 3.12. In Section 4 we study the tasking planning problem of Milnor fibrations. For instance, we show that it is possible to lower the topological complexity of work maps by restricting them in a neighbourhood of singularities (Remark 4.1). Furthermore, we compute the topological complexity of Milnor fibration and we design optimal tasking planning algorithms (Theorem 4.2).
2. Milnor fibration revisited
In [11] John Milnor showed that given a non-constant holomorphic function $g:U\subset\mathbb{C}^{n+1}\to\mathbb{C}$ with $g(0)=0$, there exists a small $\epsilon>0$ such that
(2.1) $\phi_{\epsilon}=\displaystyle\frac{g}{\|g\|}:S_{\epsilon}^{2n+1}\setminus K_{\epsilon}\to S^{1}$
is a smooth projection of a locally trivial fiber bundle (hence, it is a fibration), where
$K_{\epsilon}=g^{-1}(0)\cap S_{\epsilon}^{2n+1}$
is called the link of the singularity at origin.
Let $F_{\theta}=\phi_{\epsilon}^{-1}(e^{i\,\theta})$ be the fiber of (2.1) over a point $e^{i\,\theta}\in S^{1}$. It is a $2n$-dimensional parallelizable real manifold. In addition, using Morse theory, Milnor proved that $F_{\theta}$ has the homotopy type of a finite $CW$-complex of dimension $n$ and the link $K_{\epsilon}$ is $(n-2)$-connected, that is, the homotopy groups $\pi_{j}(K_{\epsilon})=0$, for all $j=0,\ldots,n-2.$
Let $\Sigma_{g}=\{x\in U\subset\mathbb{C}^{n+1}|\nabla g(x)=0\}$ be the set of critical points of $g$, or the singular locus of $g$. We denote just by $\Sigma$ if no mention to the function is needed.
In the case that $0\in\Sigma_{g}$ is an isolated point  [^1] In this case we say that 00 is an isolated singularity, or an isolated critical point, of $g$., Milnor gave more details about the topology of the fiber and the link. More precisely, in such a case the fiber $F_{\theta}$ has the homotopy type of a wedge of $n$-dimensional spheres $S^{n}\vee\cdots\vee S^{n},$ also known as Milnor’s bouquet of spheres, with $\mu_{g}$-copies of $S^{n}$ attached to a single common point. The number $\mu_{g}$ is called the Milnor number of $g$ at origin.
In addition, Milnor showed the existence of fiber bundle structures for real maps with isolated critical points. More precisely, Milnor considered a real analytic mapping $f:U\subset\mathbb{R}^{n}\to\mathbb{R}^{p}$ with $f(0)=0$ and $n\geq p\geq 2$ such that in some open neighborhood $U$ of the origin $0\in\mathbb{R}^{n}$, we have $\Sigma_{f}\cap U=\{0\}$, where
$\Sigma_{f}=\{x\in U|\mbox{ rank }Jf(x)\mbox{ fails to be maximal}\},$
and $Jf(x)$ denotes the Jacobian matrix of $f$ at $x$. This means that 00 is an isolated singular point of $f$.
We denote by $V=f^{-1}(0)$ and $K_{\epsilon}=V\cap S_{\epsilon}^{n-1}$, where $S_{\epsilon}^{n-1}$ is the sphere of radius $\epsilon$ centered at the origin in $\mathbb{R}^{n}$. The set $K_{\epsilon}$ is called the link of the singularity at the origin. By the transversality theorem one can prove that the diffeomorphism type of $K_{\epsilon}$ does not depend on $\epsilon$ small enough. Hence, we denote $K_{\epsilon}$ just by $K.$
Theorem 2.1.
[11, Theorem 11.2] There exists an $\epsilon_{0}>0$ small enough such that for all $\epsilon$ with $0<\epsilon\leq\epsilon_{0}$, there exists an $\eta_{0}$ with $0<\eta_{0}\ll\epsilon$ such that the restriction map
(2.2) $f_{|}:f^{-1}(S^{p-1}_{\eta})\cap B^{n}_{\epsilon}\to S^{p-1}_{\eta}$
is the projection of a smooth locally trivial fiber bundle (hence, it is a fibration) for all $\eta$ with $0<\eta\leq\eta_{0}$, where $B^{n}_{\epsilon}$ denotes the $n$-dimensional closed ball centered at the origin of radius $\epsilon$ in $\mathbb{R}^{n}$, and $S^{p-1}_{\eta}$ denotes the sphere of radius $\eta$ centered at the origin in $\mathbb{R}^{p}$. Furthermore, the fiber $F$ is a smooth compact $(n-p)$-dimensional manifold bounded by $K$.
The fibration (2.2) is known as the real Milnor fibration in the tube and its fiber $F$ as the real Milnor fiber. Rescalling the radii $\eta$ by a radial diffeomorphism we can consider the fibration over $S^{p-1},$ the unitary $(p-1)$-dimensional sphere on the target space.
An important result about the connectedness of the real Milnor fiber is given in the next proposition, which follows from [11, Lemma 11.4].
Proposition 2.2.
If the link $K=f^{-1}(0)\cap S_{\epsilon}^{n-1}$ is not empty, then the fiber $F$ of the fibration (2.2) is $(p-2)$-connected.
Moreover, we will follow the definitions and results given by D. Massey in the paper [10].
Let $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$ be a non-constant analytic map, $11\leq p\leq n-$ 1, $V:=f^{-1}(0)$ and $\Sigma_{f}$ be the set of critical points of $f.$
Let $r:\mathbb{R}^{n}\to\mathbb{R}$ be the square of the distance function to the origin, $r(x)=\|x\|^{2}$, and let $\Sigma_{(f,r)}$ be the set of critical points of the pair $(f,r)$, that is, the set of points where the gradients $\nabla r,\nabla f_{1},\ldots,\nabla f_{p}$ are linearly dependent. Note that $\Sigma_{f}\subseteq\Sigma_{(f,r)}.$ Recall that $B^{n}_{\epsilon}$ denotes the $n$-dimensional closed ball centered at the origin of radius $\epsilon$ in $\mathbb{R}^{n}$.
Definition 2.3.
[10]Given $f$ and $r$ as above.
  • (i)
    The map $f$ satisfies the Milnor condition $(a)$ at origin if $\Sigma_{f}\subset V$ in a neighborhood of the origin. Recall that $V:=f^{-1}(0)$.
  • (ii)
    The map $f$ satisfies the Milnor condition $(b)$ at origin if 00 is isolated in $V\cap\overline{\Sigma_{(f,r)}\setminus V}$ in a neighborhood of the origin, where the “bar” notation means the topological closure of the space $\Sigma_{(f,r)}\setminus V$ in $\Sigma_{(f,r)}.$
  • (iii)
    If $f$ satifies the Milnor’s condition $(a)$ (respectively $(b)$), then we say that $\epsilon_{0}>0$ is a Milnor $(a)$ radius for $f$ at origin (respectively, Milnor $(b)$ radius for $f$ at origin) provided that for all $0<\epsilon\leq\epsilon_{0}$ one gets $B_{\epsilon}^{n}\cap(\overline{\Sigma_{f}\setminus V})=\emptyset$ (respectively, $B_{\epsilon}^{n}\cap V\cap(\overline{\Sigma_{(f,r)}\setminus V})\subseteq\{0\}$).
Under condition $(iii)$ above, we simply say that $\epsilon_{0}$ is a Milnor radius for $f$ at origin, if $\epsilon_{0}$ is both a Milnor $(a)$ and Milnor $(b)$ radius for $f$ at origin.
Remark 2.4.
It can be concluded that, under Milnor’s conditions $(a)$ and $(b)$, for all regular values close to the origin of $\mathbb{R}^{p}$, the respective fibers inside the closed $\epsilon$-ball are smooth and transverse to the sphere $S_{\epsilon}^{n-1}$.
Remark 2.5.
It is easy to see that any analytic map germ $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$ with isolated critical point at origin has a Milnor’s radius. Indeed, condition (a) is satisfied because, by assumption, the critical point is isolated. Condition (b) is satisfied because $V\setminus\{0\}$ is transversal to small spheres, and so are small levels of $f$.
In a more general way, the Milnor conditions $(a)$ and $(b)$ are enough to ensure the existence of the Milnor tube fibrations as follows.
Theorem 2.6.
[10, p. 284, Theorem 4.3] Let $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$ be an analytic map germ, and $\epsilon_{0}$ be a Milnor radius for $f$ at the origin. Then, for each $0<\epsilon\leq\epsilon_{0}$, there exists $\eta$, $0<\eta\ll\epsilon$, such that
(2.3) $f_{|}:B_{\epsilon}^{n}\cap f^{-1}(B_{\eta}^{p}\setminus\{0\})\to B_{\eta}^{p}\setminus\{0\},$
is the projection of a smooth locally trivial fiber bundle.
Theorem 2.6 implies the following statement.
Corollary 2.7.
(Fibration in the Milnor tube) Let $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$ be an analytic map germ, and $\epsilon_{0}$ be a Milnor radius for $f$ at the origin. Then, for each $0<\epsilon\leq\epsilon_{0}$, there exists $\eta$, $0<\eta\ll\epsilon$, such that
(2.4) $f_{|}:B_{\epsilon}^{n}\cap f^{-1}(S_{\eta}^{p-1})\to S_{\eta}^{p-1},$
is the projection of a smooth locally trivial fiber bundle. Hence, (2.4) is a fibration.
Remark 2.8.
Let $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$ be an analytic map germ, and $\epsilon_{0}$ be a Milnor radius for $f$ at the origin. Given any commutative diagram
where $j_{0}(z)=(z,0)$, by Corollary 2.7, the map $f_{|}:B_{\epsilon}^{n}\cap f^{-1}(S_{\eta}^{p-1})\to S_{\eta}^{p-1}$ is a fibration and thus there is a continuous map $\widetilde{H}:Z\times[0,1]\to B_{\epsilon}^{n}\cap f^{-1}(S_{\eta}^{p-1})$,
such that $\widetilde{H}(z,0)=\varphi(z),\leavevmode\nobreak\ \forall z\in Z$ and $f_{|}\circ\widetilde{H}=H$. We do not know an explicit construction for such homotopy $\widetilde{H}$.
Given $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0),$ $n>p\geq 2,$ a non-constant analytic map germ and $V=f^{-1}(0),$ H. Whitney proved that one can find a stratification of $V,$ let say $S$ with finitely many subanalytics strata such that each stratum $S_{i}\in S$ that contains the origin $\{0\}$ on its closure is transversal to all sphere $S_{\epsilon}^{n-1},$ for all $\epsilon>0$ small enough. Such a stratification induces a stratification of the link $K_{\epsilon}=V\cap S_{\epsilon}^{n-1}.$ Moreover, one can use such transversality to prove that the topological type of $K_{\epsilon}$ does not depend on $\epsilon$ small enough. For this reason from now on we will denote the link only by $K.$
From [16, Theorem 8.3], we have the following statement.
Theorem 2.9.
Let $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$, $n>p\geq 2$, be an analytic map germ that satisfies Milnor’s conditions $(a)$ and $(b)$ at origin. Assume further that all cells of $K$ have dimensions $\leq n-p-1.$
  • (i)
    If $p\geq 2,$ then the total space of the Milnor fibration (2.4) is path connected.
  • (ii)
    If $p\geq 3$ then the Milnor fiber is path connected.
  • (iii)
    If $p=2$ the Milnor fiber is path connected if and only if the Milnor fibration (2.4) admits a (global) cross-section.
The next example comes from M. Ribeiro’s Ph.D thesis in [13]. It shows that for $p=2$ only the Milnor conditions $(a)$ and $(b)$ are not enough to guarantee the connectness of the Milnor tube neither the Milnor tube fibers.
Example 2.10.
Consider $f:(\mathbb{R}^{3},0)\to(\mathbb{R}^{2},0),$ $f(x,y,z)=(xy,xz).$ Note that $\Sigma_{f}=\{x=0\}$ and $V=\{x=0\}\cup\{y=z=0\},$ hence Milnor condition $(a)$ holds. Just in the same way one has $\Sigma_{(f,r)}=\{x=0\}\cup\{x^{2}=y^{2}+z^{2}\},$ $\overline{\Sigma_{(f,r)}\setminus V}=\{x^{2}=y^{2}+z^{2}\},$ and so Milnor condition $(b)$ also holds true. Then, Theorem 2.6 assures that the projection
(2.7) $f_{|}:B_{\epsilon}^{3}\cap f^{-1}(B_{\eta}^{2}\setminus\{0\})\to B_{\eta}^{2}\setminus\{0\}$
is a locally trivial fiber bundle for all $0<\eta\ll\epsilon$ small enough. Now, since the space $\{x=0\}$ disconnect the closed ball $B_{\epsilon}^{3}$ into two halves one has that the Milnor tube $B_{\epsilon}^{3}\cap f^{-1}(S_{\eta}^{1})$ is disconnected, and so are the Milnor fibers. In addition, note that the link $K$ is a CW complex of dimension $1$. On the other hand, the map $s:B_{\eta}^{2}\setminus\{0\}\to B_{\epsilon}^{3}\cap f^{-1}(B_{\eta}^{2}\setminus\{0\}),\leavevmode\nobreak\ s(a,b)=\left(\lambda,\dfrac{a}{\lambda},\dfrac{b}{\lambda}\right)$, where $\lambda=\sqrt{\dfrac{\epsilon^{2}+\sqrt{\epsilon^{4}-4\eta^{2}}}{2}}$, is a continuous cross-section of the Milnor fibration (2.7).
Furthermore, we have the following statement.
Proposition 2.11.
Let $F\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}S^{d}$ be a fibration ($d\geq 1$) with $F$ and $E$ are path connected. We have that $\pi_{d-1}(F)=0$ if and only if $\pi_{d-1}(E)=0$ and $p$ admits a global continuous cross-section.
Proof.
Consider the following piece of the long exact sequence in homotopy of the fibration $p$:
${\pi_{d}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{\#}}$${\pi_{d}(S^{d})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{d-1}(F)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{d-1}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{\#}}$${\pi_{d-1}(S^{d})=0.}$
Suppose $\pi_{d-1}(F)=0$ then $\pi_{d-1}(E)=0$ and $\pi_{d}(E)\stackrel{{\scriptstyle p_{\#}}}{{\to}}\pi_{d}(S^{d})$ is surjective and, therefore, there is a map $\lambda:S^{d}\to E$ such that $[id]=p_{\#}([\lambda])=[p\circ\lambda]$, where $[id]$ is the homotopy class of the identity map. Thus $\lambda:S^{d}\to E$ is a global continuous cross-section of $p$.
Conversely, suppose that $\pi_{d-1}(E)=0$ and the Milnor fibration admits a global continuous cross-section $\lambda:S^{d}\to E$. By the definition $[id]=[p\circ\lambda]=p_{\#}([\lambda])$ and, therefore, $p_{\#}:\pi_{d}(E)\to\pi_{d}(S^{d})$ is surjective, since $\pi_{d}(S^{d})$ is the infinity cyclic group generated by $[id]$. Thus, since
${\pi_{d}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p_{\#}}$${\pi_{d}(S^{d})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{d-1}(F)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{d-1}(E)=0.}$
is an exact sequence, we conclude that $\pi_{d-1}(F)=0$. ∎
Theorem 2.9 together with Proposition 2.11 implies the following result. It will be use in Theorem 4.2.
Proposition 2.12.
Let $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$, $n>p\geq 2$, be an analytic map germ that satisfies Milnor’s conditions $(a)$ and $(b)$ at origin. Assume further that all cells of $K$ have dimensions $\leq n-p-1.$ Recall that $F\hookrightarrow B_{\epsilon}^{n}\cap f^{-1}(S_{\eta}^{p-1})\stackrel{{\scriptstyle f_{|}}}{{\to}}S_{\eta}^{p-1}$ is the Milnor fibration (2.4). We have that $\pi_{p-2}(F)=0$ if and only if $\pi_{p-2}\left(B_{\epsilon}^{n}\cap f^{-1}(S_{\eta}^{p-1})\right)=0$ and $f_{|}$ admits a global continuous cross-section.
Milnor fibration for arrangements
We recall the Milnor fibration of arrangements from [17].
An arrangement of hyperplanes is a finite set $\mathcal{A}$ of codimension 1 linear subspaces in a finite-dimensional, complex vector space $\mathbb{C}^{d+1}$.
We will assume throughout that the arrangement is central, that is, all the hyperplanes pass through the origin. For each hyperplane $H\in\mathcal{A}$, let $L_{H}:\mathbb{C}^{d+1}\to\mathbb{C}$ be a linear form with kernel $H$.
The product
$Q:=Q(\mathcal{A})=\prod_{H\in\mathcal{H}}L_{H},$
then, is a homogeneous polynomial of degree equal to $\mid\mathcal{A}\mid$, the cardinality of the set $\mathcal{A}$. Note that $Q$ is a defining polynomial for the arrangement, unique up to a (non-zero) constant factor.
The complement of the arrangement
$M:=M(\mathcal{A})=\mathbb{C}^{d+1}-\bigcup_{H\in\mathcal{H}}H$
is a connected, smooth complex variety. Moreover, $M(\mathcal{A})$ has the homotopy type of a finite CW-complex of dimension at most $d+1$.
The polynomial map
(2.8) $Q:M(\mathcal{A})\to\mathbb{C}^{\ast}$
is the projection map of a smooth, locally trivial fiber bundle (hence, it is a fibration), known as the Milnor fibration of the arrangement. Its fiber
$F:=F(\mathcal{A})=Q^{-1}(1)$
is called the Milnor fiber of the arrangement. $F$ is a smooth, connected, orientable manifold of dimension $2d$. Moreover, $F$ has the homotopy type of a finite CW-complex of dimension $d$. If $Q$ has an isolated singularity at origin (for example, if $d=1$), then $F$ is homotopic to a wedge of $d$-spheres, and so $F$ is $(d-1)$-connected.
3. Topological complexity of a map
In this section we review the tasking planning problem of a map.
Let $PX$ denote the space of all continuous paths $\gamma:[0,1]\to X$ in a topological space $X$. Equip the path space $PX$ with the compact-open topology. From robotics (see [3]), the configuration space or state space of a mechanical system $\mathcal{S}$ is defined as the space of all possible states of $\mathcal{S}$. The robot workspace consists of all points that can be reached by the robot end-effector, i.e., the space of all tasks. The configuration space and workspace are often described as a subspace of some Euclidean space $\mathbb{R}^{n}$ and $\mathbb{R}^{p}$ respectively. The work map or kinematic map is a continuous map from the configuration space $C$ to the workspace $W$, that is, it is a continuous map
$f:C\to W$
which assigns to each state of the configuration space the position of the end-effector at that state. This map is an important object to be considered when implementing algorithms controlling the task performed by the robot manipulator.
The planning problem consists in implementing algorithms (we will say tasking planning algorithms) controlling the task performed by the robot manipulator. The input of such an algorithm (see Pavešić [12]) are pairs $(a,A)\in C\times W$ of points of the configuration space and workspace, respectively. The output for such a pair ought to be a path in the configuration space $\alpha\in PC$ such that $\alpha(0)=a$ and $f(\alpha(1))=A$.
For practical purposes, a tasking planning algorithm should depend continuously on the pair of points $(a,A)$. Indeed, if the autonomous system performs within a noisy environment, absence of continuity could lead to instability issues in the behavior of the algorithm.
Unfortunately, if a (global) continuous tasking planning algorithm of $f:C\to W$ exists implies that $f$ admits a global continuous cross-section and $W$ is contractible (see Lemma 3.8). Yet, if $f$ does not admit a global continuous cross-section or $W$ is not contractible, we could care about finding local continuous tasking planning algorithms, i.e., tasking planning algorithms $s$ defined only on a certain open set of $C\times W$, to which we refer as the domain of definition of $s$. In these terms, a tasking planner of $f$ is a set of local continuous tasking planning algorithms whose domains of definition cover $C\times W$.
The topological complexity of $f$, TC$(f)$, is then the minimal cardinality among tasking planners of $f$ (Definition 3.9), while a tasking planner of $f$ is said to be optimal if its cardinality is TC$(f)$. The design of explicit motion planners that are reasonably close to optimal is one of the challenges of modern robotics (see, for example Latombe [8] and LaValle [9]).
The concept of topological complexity of a map was introduced by Pavešić in [12], see also [19]. Here we recall the basic definitions and properties. At first we recall that, given a continuous map $f:X\to Y$, the sectional number of $f$, denoted by $\text{sec}(f)$, is the smallest number $n$ such that $Y$ can be covered by $n$ open subsets $U_{1},\ldots,U_{n}$ with the property that for every $1\leq i\leq n$ there exists a local continuous cross-section $s_{i}:U_{i}\to Y$, i.e., $f\circ s_{i}$ is equal to the inclusion map $incl_{U_{i}}:U_{i}\to Y$.
Remark 3.1.
In the case that $f$ is a fibration, we can relax the equality $f\circ s_{i}=incl_{U_{i}}$ by requiring that each of the maps $s_{i}:U_{i}\to Y$ satisfies the condition $f\circ s_{i}\simeq incl_{U_{i}}$.
For convenience, we record the following standard properties, most of which appear in Chapter 4 of [18]. If the following diagram
commutes, then $\mathrm{sec}\hskip 0.28453pt(p^{\prime})\geq\mathrm{sec}\hskip 0.28453pt(p)$. Also, for any continuous map $p:E\to B$ and any continuous map $f:B^{\prime}\to B$, note that any local continuous cross-section $s:U\to E$ of $p:E\to B$ induces a local continuous cross-section of the canonical pullback $f^{\ast}p:B^{\prime}\times_{B}E\to B^{\prime}$, called the local pullback section $f^{\ast}(s):f^{-1}(U)\to B^{\prime}\times_{B}E$, simply by defining
$f^{\ast}(s)(b^{\prime})=\left(b^{\prime},\left(s\circ f\right)(b^{\prime})\right).$
Thus,
(3.11) $\text{sec}(f^{\ast}p)$ $\leq\text{sec}(p).$
From [20, p. 1619], a quasi pullback means a strictly commutative diagram
such that, for any strictly commutative diagram as the one on the left hand-side of (3.17), there exists a (not necessarily unique) map $h:Z\to X^{\prime}$ that renders a strictly commutative diagram as the one on the right hand-side of (3.17).
Note that such a condition amounts to saying that $X^{\prime}$ contains the canonical pullback $Y^{\prime}\times_{Y}X$ determined by $f$ and $\varphi$ as a retract in a way that is compatible with the mappings into $X$ and $Y^{\prime}$.
The proof of the following statement is very technical and will be use in Proposition 3.12 and Theorem 4.2.
Lemma 3.2.
Let $p:E\to B$ be a continuous map. If the following square
[…]
Proof.
Since $p^{\prime}$ is a quasi pullback, we have the following commutative triangle
and thus $\text{sec}(f^{\ast}p)\geq\text{sec}(p^{\prime})$. Similarly, since $f^{\ast}p$ is the canonical pullback, we have the following commutative triangle
and thus $\text{sec}(p^{\prime})\geq\text{sec}(f^{\ast}p)$. Hence, the equality $\text{sec}(p^{\prime})=\text{sec}(f^{\ast}p)$ holds. By the inequality (3.11), we obtain $\mathrm{sec}\hskip 0.28453pt(p^{\prime})\leq\mathrm{sec}\hskip 0.28453pt(p)$.
The Lusternik-Schnirelmann category (LS category) or category of a topological space $X$ (see [4]), denoted by cat$(X)$, is the least integer $n$ such that $X$ can be covered with $n$ open sets which are all contractible within $X$. We have $\text{cat}(X)=1$ iff $X$ is contractible. The LS category is a homotopy invariant, i.e., if $X$ is homotopy equivalent to $Y$ then $\text{cat}(X)=\text{cat}(Y)$. Note that $\text{cat}(S^{n})=2$ for any $n\geq 0$.
The sectional number satisfies the following basic properties.
Lemma 3.3.
Let $p:E\to B$ be a map and $h^{\ast}$ be any multiplicative cohomology theory. Then
  1. (1)
[…]
  2. (2)
    [18, Proposição 4.3.17-(3), pg. 138] If there are $x_{1},\ldots,x_{k}\in h^{\ast}(B)$ with
    $p^{\ast}(x_{1})=\cdots=p^{\ast}(x_{k})=0\text{ and }x_{1}\cup\cdots\cup x_{k}\neq 0,$
    then
[…]
    Where $p^{\ast}:h^{\ast}(B)\to h^{\ast}(E)$ is the induced homomorphism in cohomology.
Remark 3.4.
In this paper we will use Item (2) of Lemma 3.3 for $h^{\ast}$ as being the singular cohomology $H^{\ast}(-;R)$ with any coefficient ring $R$ (as was presented by James in [7, pg. 342]).
Item (iii) from Theorem 2.9 implies a characterisation of the connectivity of the Milnor fiber in terms of the sectional number.
Proposition 3.5.
[…]
Proof.
Suppose that $\text{sec}(f_{|})=2$. Then $f_{|}$ does not admit a continuous (global) cross-section, and by Item (iii) from Theorem 2.9, we have that the Milnor fiber is not path connected.
Now, suppose that the Milnor fiber is not path connected. Then, by Item (iii) from Theorem 2.9, we obtain that $f_{|}$ does not admit a continuous (global) cross-section and thus $\text{sec}(f_{|})\geq 2$. Furthermore, by Item (1) of Lemma 3.3, we have that $\text{sec}(f_{|})\leq\text{cat}(S^{1}_{\eta})$. Recall that $\text{cat}(S^{1}_{\eta})=2$ and hence $\text{sec}(f_{|})=2$. ∎
Lemma 3.3 implies the following example.
Example 3.6.
[…]
A direct application of Proposition 3.5 together with Example 3.6 is the following statement.
Proposition 3.7.
[…]
For a topological space $X$, recall that $PX$ denote the space of all paths in $X$ with the compact-open topology. Consider the evaluation fibration
$e_{0,1}:PX\to X\times X,\leavevmode\nobreak\ e_{0,1}(\gamma)=\left(\gamma(0),\gamma(1)\right).$
For any continuous map $f:X\to Y$, consider the composite
$e_{f}:PX\to X\times Y,\leavevmode\nobreak\ e_{f}=(1_{X}\times f)\circ e_{0,1}.$
A tasking planning algorithm is a cross-section $s\colon X\times Y\to PX$ of the map $e_{f}$, i.e.,  a (not necessarily continuous) map satisfying $e_{f}\circ s=1_{X\times Y}$. Note that, by definition, if there is a tasking planning algorithm of $f$ then $f$ is surjective. Furthermore, we have the following statement.
Lemma 3.8.
Let $f:X\to Y$ be any continuous map. If a continuous tasking planning algorithm of $f$ exists then $f$ admits a global continuous cross-section and the space $Y$ is contractible.
Proof.
Suppose that $\sigma:X\times Y\to PX$ is a continuous cross-section of $f$, that is, $\sigma$ is a continuous map and for any $(x,y)\in X\times Y$, we have that $\sigma(x,y)(0)=x$ and $f(\sigma(x,y)(1))=y$. Set $x_{0}\in X$ and consider the map $s:Y\to X,\leavevmode\nobreak\ s(y)=\sigma(x_{0},y)(1)$. Note that $s$ is a global continuous cross-section of $f$. In addition, the homotopy $H:Y\times[0,1]\to Y$, given by $H(y,t)=f(\sigma(x_{0},y)(t))$, satisfies $H(y,0)=f(x_{0})$ and $H(y,1)=y$, for any $y\in Y$. Hence, $Y$ is contractible.
Lemma 3.8 forces the following definition.
Definition 3.9.
The topological complexity TC$(f)$ of a continuous map $f:X\to Y$ is the sectional number of the map $e_{f}$. In other words the topological complexity of $f$ is the smallest positive integer TC$(f)=k$ for which the product $X\times Y$ is covered by $k$ open subsets $X\times Y=U_{1}\cup\cdots\cup U_{k}$ such that for any $i=1,2,\ldots,k$ there exists a continuous cross-section $s_{i}:U_{i}\to PX$ of $e_{f}$ over $U_{i}$ (i.e., $e_{f}\circ s_{i}=incl_{U_{i}}$). If no such $k$ exists we will set TC$(f)=\infty$.
Note that, any collection $s=\{s_{i}:U_{i}\to PX\}_{i=1}^{k}$, where $\{U_{i}\}_{i=1}^{k}$ is an open covering of $X\times Y$ and each $s_{i}:U_{i}\to PX$ is a continuous local cross-section of $e_{f}$ over $U_{i}$; implies a tasking planning algorithm $s:X\times Y\to PX$ given by $s(x,y)=s_{i}(x,y)$ where $i$ is the minimal index in such that $(x,y)\in U_{i}$. Any such tasking planning algorithm $s:=\{s_{i}:U_{i}\to PX\}_{i=1}^{k}$ is called optimal if $k=\text{TC}(f)$.
For convenience, we record the following standard properties of topological complexity (see [19]):
Remark 3.10.
  1. (1)
[…]
  2. (2)
    Consider the diagram of maps $E^{\prime}\stackrel{{\scriptstyle p^{\prime}}}{{\to}}E\stackrel{{\scriptstyle p}}{{\to}}B\stackrel{{\scriptstyle p^{\prime\prime}}}{{\to}}B^{\prime}$. If $p$ admits a continuous cross-section, then
    • $a$)
[…]
    • $b$)
[…]
The topological complexity of the identity map $1_{X}:X\to X$, $\text{TC}(1_{X})=\text{TC}(X)$, coincides with the topological complexity (a la Farber) of $X$ (see Farber [6]). We have $\text{TC}(X)=1$ iff $X$ is contractible. The TC is also a homotopy invariant, i.e., if $X\simeq Y$ then $\text{TC}(X)=\text{TC}(Y)$.
Example 3.11.
[Motion planning over spheres] By Farber [6], the TC for spheres is as follows:
[…]
Furthermore, suppose $m$ is odd. Let $v$ denote a fixed continuous unitary tangent vector field on $S^{m}$, say $v(x_{1},y_{1},\ldots,x_{\ell},y_{\ell})=(-y_{1},x_{1},\ldots,-y_{\ell},x_{\ell})$ with $m+1=2\ell$.
A planning algorithm to $S^{m}$ with $m$ odd is given by $s:=\{s_{i}:U_{i}\to PS^{m}\}_{i=1}^{2}$, where
$U_{1}$ $=$ $\{(\theta_{1},\theta_{2})\in S^{m}\times S^{m}\mid\leavevmode\nobreak\ \leavevmode\nobreak\ \theta_{1}\neq-\theta_{2}\},$
$U_{2}$ $=$ $\{(\theta_{1},\theta_{2})\in S^{m}\times S^{m}\mid\leavevmode\nobreak\ \leavevmode\nobreak\ \theta_{1}\neq\theta_{2}\},$
with local algorithms:
$s_{1}(\theta_{1},\theta_{2})(t)=\dfrac{(1-t)\theta_{1}+t\theta_{2}}{\parallel(1-t)\theta_{1}+t\theta_{2}\parallel}\text{ for all }(\theta_{1},\theta_{2})\in U_{1}.$
Before to define $s_{2}$, we can consider the subset $F:=\{(\theta_{1},\theta_{2})\in S^{m}\times S^{m}\mid\leavevmode\nobreak\ \leavevmode\nobreak\ \theta_{1}=-\theta_{2}\}$ and for $(\theta_{1},\theta_{2})\in F$ define
$\alpha(\theta_{1},\theta_{2})(t)=\begin{cases}s_{1}(\theta_{1},v(\theta_{1}))(2t),&\hbox{ $0\leq t\leq\dfrac{1}{2}$;}\\ s_{1}(v(\theta_{1}),\theta_{2})(2t-1),&\hbox{ $\dfrac{1}{2}\leq t\leq 1$.}\end{cases}$
Now, for all $(\theta_{1},\theta_{2})\in U_{2}$, set
$s_{2}(\theta_{1},\theta_{2})(t)=\begin{cases}s_{1}(\theta_{1},-\theta_{2})(2t),&\hbox{ $0\leq t\leq\dfrac{1}{2}$;}\\ \alpha(-\theta_{2},\theta_{2})(2t-1),&\hbox{ $\dfrac{1}{2}\leq t\leq 1$,}\end{cases}$
Moreover, we present an optimal algorithm for even-dimensional spheres. Suppose $m$ is even. Let $\nu:S^{m}\to\mathbb{R}^{m+1}$ denote a fixed continuous tangent vector field on $S^{m}$, which vanishes at two points $-1,1\in S^{m}$ and is nonzero for any $x\in S^{m}-\{-1,1\}$. Indeed, we can take
$\nu(x_{1},x_{2},x_{3},\ldots,x_{m},x_{m+1})=(0,-x_{3},x_{2},\ldots,-x_{m+1},x_{m}).$
Note that $\nu(1,0,\ldots,0)=0$, $\nu(-1,0\ldots,0)=0$ and $\nu(x)\neq 0$ for any $x\in S^{m}-\{-1,1\}$.
A planning algorithm to $S^{m}$ with $m$ even is given by $\kappa:=\{\kappa_{i}:V_{i}\to PS^{m}\}_{i=1}^{3}$, where
$V_{1}$ $=$ $\{(\theta_{1},\theta_{2})\in S^{m}\times S^{m}\mid\leavevmode\nobreak\ \leavevmode\nobreak\ \theta_{1}\neq-\theta_{2}\},$
$V_{2}$ $=$ $\{(\theta_{1},\theta_{2})\in S^{m}\times S^{m}\mid\leavevmode\nobreak\ \leavevmode\nobreak\ \theta_{1}\neq\theta_{2}\text{ and }\theta_{2}\neq-1,1\},$
with local algorithms:
$\kappa_{1}(\theta_{1},\theta_{2})(t)=\dfrac{(1-t)\theta_{1}+t\theta_{2}}{\parallel(1-t)\theta_{1}+t\theta_{2}\parallel}\text{ for all }(\theta_{1},\theta_{2})\in V_{1}.$
Before to define $\kappa_{2}$, we can consider the subset $G:=\{(\theta_{1},\theta_{2})\in S^{m}\times S^{m}\mid\leavevmode\nobreak\ \leavevmode\nobreak\ \theta_{1}=-\theta_{2}\text{ and }\theta_{2}\neq-1,1\}$ and for $(\theta_{1},\theta_{2})\in G$ define
$\beta(\theta_{1},\theta_{2})(t)=\begin{cases}\kappa_{1}(\theta_{1},\nu(\theta_{1}))(2t),&\hbox{ $0\leq t\leq\dfrac{1}{2}$;}\\ \kappa_{1}(\nu(\theta_{1}),\theta_{2})(2t-1),&\hbox{ $\dfrac{1}{2}\leq t\leq 1$.}\end{cases}$
Now, for all $(\theta_{1},\theta_{2})\in V_{2}$, set
$\kappa_{2}(\theta_{1},\theta_{2})(t)=\begin{cases}\kappa_{1}(\theta_{1},-\theta_{2})(2t),&\hbox{ $0\leq t\leq\dfrac{1}{2}$;}\\ \beta(-\theta_{2},\theta_{2})(2t-1),&\hbox{ $\dfrac{1}{2}\leq t\leq 1$,}\end{cases}$
Now, $V_{1}\cup V_{2}$ covers everything except the points $(-1,1)$ and $(1,-1)$. We recall the stereographic projection with respect to the north pole $p_{N}=(0,\ldots,0,1)$:
$p:S^{m}-\{p_{N}\}\to\mathbb{R}^{m},\leavevmode\nobreak\ (x_{1},\ldots,x_{m+1})=\left(\dfrac{x_{1}}{1-x_{m+1}},\ldots,\dfrac{x_{m}}{1-x_{m+1}}\right)$
whose inverse is given by
$q:\mathbb{R}^{m}\to S^{m}-\{p_{N}\},\leavevmode\nobreak\ y=(y_{1},\ldots,y_{m})=\left(\dfrac{2y_{1}}{\parallel y\parallel^{2}+1},\ldots,\dfrac{2y_{m}}{\parallel y\parallel^{2}+1},\dfrac{\parallel y\parallel^{2}-1}{\parallel y\parallel^{2}+1}\right).$
Set $Y=S^{m}-\{p_{N}\}$. This means that we may take $V_{3}=Y\times Y$. The algorithm $\kappa_{3}:V_{3}\to PS^{m}$ is given by
$\kappa(\theta_{1},\theta_{2})(t)=q\left((1-t)p(\theta_{1})+tp(\theta_{2})\right)\text{ for }(\theta_{1},\theta_{2})\in V_{3}\text{ and }t\in[0,1].$
The proof of the following proposition is also fundamental for our purposes.
Proposition 3.12.
If $p:E\to B$ is a fibration then
[…]
Proof.
We will check $\text{TC}(p^{\prime}\circ p)\leq\text{TC}(p^{\prime})$ for any map $p^{\prime}:B\to B^{\prime}$. In particular, $\text{TC}(p)\leq\text{TC}(B)$. By Lemma 3.2, it is sufficient to prove that the following diagram:
is a quasi pullback. Indeed, for any $\beta:X\to PB$ and any $\alpha:X\to E\times B^{\prime}$ satisfying $e_{p^{\prime}}\circ\beta=(p\times 1_{B^{\prime}})\circ\alpha$, we need to see that there exists $H:X\to PE$ such that $e_{p^{\prime}\circ p}\circ H=\alpha$ and $p_{\#}\circ H=\beta$.
Note that we have the following commutative diagram:
where $p_{1}$ is the projection onto the first coordinate. Because $p$ is a fibration, there exists $H:X\times I\to E$ satisfying $H\circ i_{0}=p_{1}\circ\alpha$ and $p\circ H=\beta$, thus we obtain the desired quasi pullback.
The second part of the proposition follows from Item (1) of Remark 3.10.
4. Topological complexity of Milnor fibration
Now, we are ready to study the tasking planning problem of Milnor fibrations. Explicitly, we compute the topological complexity of Milnor fibration and we design optimal tasking planning algorithms (Theorem 4.2). By the proof of Theorem 4.2 together with Remark 2.8 this algorithms strongly depends of the Milnor fibration.
We consider $n>p\geq 2$ and $f:(\mathbb{R}^{n},0)\longrightarrow(\mathbb{R}^{p},0)$ be an analytic map germ which satisfies the Milnor’s conditions $(a)$ and $(b)$. In particular, we consider the Milnor fibration as work map:
(4.24) $f_{\mid}:f^{-1}(S^{p-1}_{\eta})\cap B^{n}_{\epsilon}\to S^{p-1}_{\eta},$
where $0<\eta\ll\epsilon\leq\epsilon_{0},$ and $\epsilon_{0}$ is a Milnor’s radius for $f$ at origin. Recall that, $E=E(\eta,\epsilon)=f^{-1}(S^{p-1}_{\eta})\cap B^{n}_{\epsilon}$ denotes the Milnor tube.
Remark 4.1.
  1. (1)
    By Item (1) of Lemma 3.3 together with Item (1) of Remark 3.10, Proposition 3.12 and Example 3.11, we obtain
[…]
  2. (2)
    By the proof of Proposition 3.12, any optimal algorithm from $S^{p-1}$ induces an algorithm to $f_{\mid}$ not necessarily optimal. However, it is optimal when $f_{\mid}$ admits a continuous cross-section or $p$ is even.
Now, we compute the value of $\text{TC}(f_{\mid})$ when
  1. (1)
    $f_{\mid}:M(\mathcal{A})\to\mathbb{C}^{\ast}$ is a Milnor fibration of a central arrangement $\mathcal{A}$.
  2. (2)
    $f:(\mathbb{C}^{n},0)\to(\mathbb{C},0)$ is a holomorphic function germ (not necessarily with $n\geq 2+\text{dim}_{\mathbb{C}}\Sigma_{f}$).
  3. (3)
    $f:(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$, $n>p\geq 2$, is an analytic function germ with $p$ even (not necessarily with an isolated singular point at origin).
  4. (4)
    $f:(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$, $n>p\geq 2$, is an analytic function germ with $p$ odd, with an isolated singular point at origin and not empty link $K=f^{-1}(0)\cap S_{\epsilon}^{n-1}$.
  5. (5)
    $f:(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$, $n>p\geq 2$, is an analytic function germ with $p$ odd, with all cells of $K$ have dimensions $\leq n-p-1$ and $\pi_{p-2}(F)=0$.
Theorem 4.2 (Principal Theorem).
We have
[…]
Furthermore, we design optimal tasking planning algorithms with $2$ and $3$ regions of continuity, respectively.
Proof.
The cases (1), (2) and (3) follows directly from Remark 4.1. The case (4) and (5) is more difficult, however, by the Milnor fibration theory we have that $f_{\mid}$ admits a global continuous cross-section whenever the link $K=f^{-1}(0)\cap S_{\epsilon}^{n-1}$ is not empty, see [11, p. 101] and Proposition 2.12, respectively. Then we can apply again Proposition 3.12 and thus Example 3.11 implies the result.
Now, we present optimal tasking planning algorithms with $2$ and $3$ regions of continuity for the Milnor fibration $f_{\mid}:E(\eta,\epsilon)\to S^{p-1}$, respectively. Recall the following quasi pullbacks:
where $B^{\prime}=E(\eta,\epsilon)\times S^{p-1}$, $B=S^{p-1}\times S^{p-1}$ and $E=PS^{p-1}$. Recall that $H$ is given by the following commutative diagram (see Remark 2.8):
where $X=B^{\prime}\times_{B}E$.
For $p$ even From Proposition 3.12, the optimal algorithm $s:=\{s_{i}:U_{i}\to PS^{p-1}\}_{i=1}^{2}$ on the sphere $S^{p-1}$ (see Example 3.11) induces an optimal tasking algorithm for $f_{\mid}$ with 2 local algorithms, say $\hat{s}:=\{\hat{s}_{i}:\widehat{V}_{i}\to PE(\eta,\epsilon)\}_{i=1}^{2}$, where $\widehat{V}_{i}=(f_{\mid}\times 1)^{-1}(U_{i})\subset E(\eta,\epsilon)\times S^{p-1}$ and $\hat{s}_{i}(v)=H(v,s_{i}\circ(f_{\mid}\times 1)(v))$.
For $p$ odd and with an isolated singular point at origin Again, from Proposition 3.12, the optimal algorithm $\kappa:=\{\kappa_{i}:V_{i}\to PS^{p-1}\}_{i=1}^{3}$ on the sphere $S^{p-1}$ (see Example 3.11) induces an optimal tasking algorithm for $f_{\mid}$ with 3 local algorithms, say $\hat{\kappa}:=\{\hat{\kappa}_{i}:\widehat{V}_{i}\to PE(\eta,\epsilon)\}_{i=1}^{3}$, where $\widehat{V}_{i}=(f_{\mid}\times 1)^{-1}(V_{i})\subset E(\eta,\epsilon)\times S^{p-1}$ and $\hat{s}_{i}(v)=H(v,\kappa_{i}\circ(f_{\mid}\times 1)(v))$. ∎
By the proof of Theorem 4.2 together with Remark 2.8 this algorithms strongly depends of the Milnor fibration. Hence, we stay the following question.
Question: Let $f=(f_{1},\ldots,f_{p}):(\mathbb{R}^{n},0)\to(\mathbb{R}^{p},0)$ be an analytic map germ, and $\epsilon_{0}$ be a Milnor radius for $f$ at the origin. Given any commutative diagram
where $j_{0}(z)=(z,0)$. How to construct a continuous map $\widetilde{H}:Z\times[0,1]\to B_{\epsilon}^{n}\cap f^{-1}(S_{\eta}^{p-1})$,
such that $\widetilde{H}(z,0)=\varphi(z),\leavevmode\nobreak\ \forall z\in Z$ and $f_{|}\circ\widetilde{H}=H$?
This of course suits best for future implementation-oriented objectives.
Conflict of interest
The authors declare that they have no conflict of interest.